Combustion of hydrocarbons such as ethane ( C 2 H 6 ) produces carbon dioxide, a "greenhouse gas." Greenhouse gases in the Earth's atmosphere can trap the Sun's heat, raising the average temperature of the Earth. For this reason there has been a great deal of international discussion about whether to regulate the production of carbon dioxide. 1. Write a balanced chemical equation, including physical state symbols, for the combustion of gaseous ethane into gaseous carbon dioxide and gaseous water. 2. Suppose 0.330 kg of ethane are burned in air at a pressure of exactly 1 atm and a temperature of 19.0 °C . Calculate the volume of carbon dioxide gas that is produced. Be sure your answer has the correct number of significant digits.
Answer #1

2C2H6(g) + 7O2(g) --------------> 4CO2(g) + 6H2O(g)

2 moles of C2H6 combustion with excess of O2 to gives 4 moles of CO2

2*30g of C2H6 combustion with excess of O2 to gives 4 moles of CO2

330g of C2H6 combustion with excess of O2 to gives 4mole*330g/(2*30g) = 22 moles of CO2

n = 22 moles

T = 19+273   = 292K

P =1 atm

PV = nRT

V   = nRT/P

      = 22*0.0821*292/1   = 527.41L

The volume of CO2 = 527.41L >>>>answer